OpenStudy (anonymous):

What is the geometric mean of 3 and 21? What is the geometric mean of 3 and 21? @Mathematics

hero (hero):

Geometry isn't mean

OpenStudy (anonymous):

its a question on my homework that's what it says

OpenStudy (mathmagician):

it is sqrt(3*21)=sqrt(63)

OpenStudy (anonymous):

What the geometric mean, means is that we will take the square root of the product of the two numbers you where given

OpenStudy (anonymous):

thank you <3

hero (hero):

It would have been easier if they had just said take the square root of the product of 3 and 21

OpenStudy (mathmagician):

you're welcome

OpenStudy (anonymous):

\[\sqrt{3*21}=3 \sqrt{7} \]

OpenStudy (anonymous):

From google: geometric mean "Mathematical definition: The nth root of the product of n numbers."